OpenStudy (horotat):

How To Solve Integral Of : (1+x)/(1-2x-(v^2))

OpenStudy (horotat):

\[\int\limits_{}^{} \frac{ 1+x }{ 1-2\times x - x^2}\] = ?

sam (.sam.):

\[\int\limits \frac{1+x}{1-2 x-x^2} \, dx\]

sam (.sam.):

Try u-sub u=1-2x-x^2 \[du=-2-2x dx \\ \\ du=-2(1+x)dx \\ \\ -\frac{du}{2} \\ \\ -\frac{1}{2}\int\limits \frac{du}{u}\]

sam (.sam.):

integral of 1/u is ln(u) \[-\frac{1}{2}(\ln(u))+c \\ \\ -\frac{1}{2}\ln(1-2x-x^2)+c\] Got it?

OpenStudy (horotat):

Good Job Dude :)